Steam enters a turbine 80 lbm of steam per minute with an enthalpy of 1700 Btu/lb and velocity of 98 ft/s. It leaves the turbine at 1400 Btu/lb at 880 ft/s. The heat loss is 90,000 Btu/hr. Find the horsepower output of turbine. 


W = ____ HP


Answer in 2 decimal places


1
Expert's answer
2021-08-25T09:07:59-0400

Energy velence steam enter a turbine

"E_{in}=E_{out}"



"K E_1+P E_1+H_1=KE_2+PE_2+H_i2+Q+W\\""K E_1+P E_1+m_sh_1=KE_2+PE_2+m_sh_2+Q+W\\""P=\frac{1}{2}m(v_1^2-v_2^2)+m_s(h_1-h_2)-Q"

Put value


"P=\frac{1}{2}\times\frac{80}{60}\times\frac{1}{32.2\times550}(880^2-98^2)+\frac{80}{42.4}(1700-1400)-\frac{90000}{2545}"

"P=28.75+566.03-35.37=560.41hp"